# RAG-Enhanced Commit Message Generation

### v2

@@ Abstract.
Commit message is one of the most important textual information in software development and maintenance. However, it is time-consuming and labor-intensive to write commit messages manually. Commit Message Generation (CMG) has become a research hotspot in automated software engineering. Researchers have proposed several methods for CMG and achieved great results. In recent years, CodeBERT, CodeT5, and other Pre-trained Language Models (PLMs) for code have been proposed. These models can be easily transferred to code-related downstream tasks including CMG with simple fine-tuning and can achieve impressive performance. Moreover, Large Language Models (LLMs) with code capabilities (e.g., ChatGPT, Llama 3, Gemma) can be directly applied to various tasks by designing instruct prompts without training. This brings new possibilities to the CMG task. Prior studies have preliminarily explored the utilization of retrieval techniques for CMG tasks but were limited to relying on a single specific model for generation. It remains unclear what effects would emerge from combining advanced retrieval techniques with various generation models. In this work, we propose REACT, a novel REtrieval-Augmented framework for CommiT message generation, which effectively integrates advanced retrieval techniques with different PLMs and LLMs and can broadly enhance the performance of various models on the CMG task. Specifically, we design and build a hybrid retriever to retrieve the most relevant code diff and commit message pair from the code base as an “exemplar”. Then, the retrieved pair is utilized to guide and enhance the generation of commit messages by PLMs and LLMs through fine-tuning and in-context learning. Our approach is evaluated on a widely-used dataset. The experimental results show that REACT significantly enhances the performance of various models on the CMG task, improving the BLEU score of CodeT5 by up to 55%, boosting Llama 3’s BLEU score by 102%, and substantially surpassing all baselines, achieving a new SOTA. This demonstrates the effectiveness and broad applicability of our framework that can enhance CMG by a large margin.
@@ 1. Introduction
In this context, Commit Message Generation (CMG) has become a hot topic in automated software engineering and has garnered attention from this research community. CMG task aims to take the differences between two versions of code as input, typically in the form of a code diff file .diff generated by Git, aka code change, and then generates the corresponding commit message. Initially, rule-based approaches were used in CMG (Shen et al., 2016; Linares-Vásquez et al., 2015; Buse and Weimer, 2010), where predefined rules or templates were utilized for generation. Some retrieval-based approaches leverage information retrieval (IR) techniques to suggest commit messages from similar code diff (Liu et al., 2018; Hoang et al., 2020). With the rapid development of deep learning, many learning-based methods have recently emerged (Jiang et al., 2017; Loyola et al., 2017; Xu et al., 2019; Dong et al., 2022; Liu et al., 2019). They treat CMG as a neural machine translation task, taking code diff as the input of neural network model to generate commit messages as output. There are also hybrid approaches (Liu et al., 2022; Wang et al., 2022; Shi et al., 2022) that leverage both IR and deep learning techniques.
More recently, various language models possessing code capabilities have been proposed, including code-specific pre-trained language models such as CodeT5 (Wang et al., 2021) and general large language models, notably ChatGPT [^1] https://chatgpt.com/, which have attracted the vast attention of the research community. In this paper, we refer to these code-capable language models as Code Language Models (CLMs). Different from traditional models, CLMs can be adapted for downstream code-related tasks by merely fine-tuning or prompting (Wei et al., 2021), rather than training a model from initial weights. Such approaches not only conserve resources and training time but also can achieve better performance. Researchers have explored applying them to the CMG task (Zhang et al., 2024c, a), showing promising prospects.
@@
In this paper, we propose REACT as a novel REtrieval-Augmented framework for CommiT message generation. REACT comprises three phases: Retrieve, Augment, and Generate to generate commit messages for given code diff. Retrieve: We design a hybrid retriever to retrieve the most relevant diff and its corresponding commit message from a comprehensive and high-quality source database. The retrieved diff-message pair will be used to guide the subsequent generation of commit message. Augment: The query diff and retrieved diff-message pair need to be combined as the input of the models, that is, input augmentation. The augmenter concatenates them with pre-defined special tokens or fills them into a prompt template before passing the result to the generator for the next phase. Generate: The generator receives augmented inputs and generates a commit message corresponding to the query diff under the guidance of the retrieved pair. We use various CLMs as the backbone of the generator instead of training a model from scratch. This can fully leverage the prior knowledge in CLMs and generate messages more effectively. Some generators with PLM backbones require training to learn how to leverage the additional information provided by the retrieved pair to enhance commit message generation. If LLM is used as the generator, enhancement can be achieved through in-context learning (Dong et al., 2023) by embedding the retrieved pairs into the input prompt, eliminating the need for training. The final commit message generated through three stages is enhanced by the additional information or similar patterns provided by the exemplar.
@@ Results.
We conducted comprehensive experiments based on a widely-used CMG dataset. Before assessing the enhancement effects of REACT, we first reported the performance of directly using CLMs to generate commit messages. The experimental results showed that without integrating the REACT framework, directly generating commit messages with PLMs yielded remarkable results, surpassing all baselines comprehensively. Specifically, CodeT5’s BLEU score was 21% higher than the best baseline method, demonstrating the superiority of PLMs. Although LLMs showed mediocre scores on some metrics, Llama 3’s METEOR score still surpassed all baselines without additional training. When integrating CLMs into REACT, the model’s performance had broad and significant improvements. CodeT5’s BLEU score increased by 55% compared to direct application, and Llama 3’s BLEU score increased by 102%. The results established the effectiveness and wide applicability of the REACT framework, setting a new SOTA for the CMG task.
@@
  • A novel retrieval-augmented framework for CMG. We propose REACT, a novel retrieval-augmented framework for enhancing commit message generation. REACT can effectively integrate advanced retrieval techniques with various models for better CMG.
  • Extensive experiments are designed to evaluate the performance of various CLMs on the CMG task. To the best of our knowledge, this paper is the first to comprehensively evaluate the performance of various CLMs on the CMG task, including PLMs (e.g., CodeT5, UniXCoder, etc.) and LLMs (e.g., ChatGPT, Llama 3, Gemma). Experimental results show that, benefiting from the code understanding ability acquired during pre-training, CLMs achieved impressive performance on CMG with simple fine-tuning or prompting with basic instruction.
@@
  • The replication package is available online (Zhang et al., 2024b).
@@ 3. Methodology
This section presents the methodology of our proposed three-phrase framework, REACT. We first provide a brief overview of REACT. Then, the three phases of REACT are addressed in detail.
@@ 3.1. Overview
As shown in Fig. 2, REACT comprises three phases, i.e., Retrieve, Augment, and Generate to generate the commit message for a specified code diff. In this paper, we refer to the code diff that awaits the generation of its corresponding commit message as the “query diff”. Firstly, REACT retrieves the most relevant diff-message pair from the source database using a hybrid retriever. Then, the query diff and the retrieved diff-message pair are sent together to the augmenter for augmentation. Finally, the augmented input is fed into the generator, which generates a commit message under the guidance of the retrieved pair.
@@ 3.2. Phase I: Retrieve
The goal of this phase is to retrieve the most relevant diff-message pair for guidance. which is done by the following two components:
@@ 3.2.1. Hybrid Retriever
To more accurately assess the similarity between two code diffs and retrieve the most relevant diff-message pair based on that similarity, we designed an advanced hybrid retriever that involves combining two similarity scores with weighted fusion. BM25 (Robertson and Zaragoza, 2009) is a relevance scoring algorithm commonly used in information retrieval and search engines. It measures the relevance of a document to a given query by considering factors such as term frequency, document length, and term saturation. BM25 treats query diff as a bag-of-words representation and computes lexical similarity scores between query diff and each of the candidates. We also employ a transformer encoder-based similarity scoring method. CodeT5+ (Wang et al., 2023a), an improved version of CodeT5 (Wang et al., 2021), is a cutting-edge PLM for code with the architecture of transformer encoder-decoder. Through its pre-training process, it learned rich representations from code data. We extract the encoder component of CodeT5+ and then utilize it to transform the code diff text into a dense 256-dimensional vector embedding which is generally considered to encapsulate the semantic information of the code diff. By calculating the cosine similarity score between two vectors, we can obtain the similarity score between two code diffs. Following the normalization of the scores to a common scale, we combine the scores obtained from these two methods with equal weights (1:1) in our experiment, and we refer to it as the hybrid score. After calculating the hybrid scores between the query diff and all diffs in the source database, the diff-message pair with the highest score is retrieved as the most relevant pair. The hybrid retriever is expected to be more robust and effective compared to single retrievers that rely on only one similarity scoring method.
@@ 3.2.2. Source Database
To ensure that each query diff can retrieve a relevant diff-message pair as an exemplar and that the retrieved exemplar is of high quality, containing sufficient information to guide generation, a comprehensive, high-quality source database is crucial. To avoid reinventing the wheel, we reuse the data from CommitBench (Schall et al., 2024), a recently proposed dataset, to construct the source database used in our work. CommitBench collects commits from 72,000 publicly available GitHub repositories, which is the largest collection scope among existing datasets, ensuring the comprehensiveness of the dataset. Regarding the quality of this dataset, firstly, the repositories collected by CommitBench are non-fork repositories that have been used by at least one other project and have a reasonable number of stars. Additionally, they performed an overall filtering process, designing ten filtering rules (e.g. bot messages, messages less than 8 in length, etc.) to exclude worthless messages. The resulting source database contains over 1.6 million diff-message pairs. During the preprocessing stage, we persisted the 256-dimensional vector embedding of code diffs, obtained through the encoder in the hybrid retriever, as a new column in the source database to avoid additional computational overhead.
@@ 3.3. Phase II: Augment
This phase is responsible for combining the query diff and the retrieved pair into specific formats to augment the inputs. For convenience, We denote <query diff>, <retrieved diff>, <retrieved msg> as the query diff for generation, the retrieved code diff, and the corresponding retrieved commit message, respectively. Considering the differences between PLMs and LLMs, we design the following two input augmentation methods for both:
@@ 3.3.1. For PLMs
PLMs require a specific input format to serve as the generator in REACT. To clearly distinguish between different input components, we define three special tokens: [QUERY], [DIFF], and [MSG]. The augmenter concatenates these special tokens with query diff and retrieved pair into the following form:
@@ 3.3.2. For LLMs
We manually construct a prompt template for LLMs as Figure 3 shown, enabling them to generate commit messages by referencing the retrieved exemplar. The augmenter fills in the corresponding content into the prompt template and then inputs it into LLMs for generation.
@@ 3.4. Phase III: Generate
The generator receives the augmented input from phase II and generates the final commit message under the guidance of the retrieved pair inserted in the input. In this paper, we adopt various different models as the generator, including PLMs and LLMs, and conduct a comprehensive study. The specific model selection will be detailed in Section 4.3. In this subsection, we discuss the two types of generators used in REACT and what adjustments they need to make to perform generating.
@@ 3.4.2. LLMs
In contrast to PLMs, LLMs can serve as the generator to generate commit messages without training. It benefits from the powerful generalization ability of LLMs, which can complete the specified tasks with only prompting (Wei et al., 2021). Meanwhile, through proper instruction designing as shown in Figure 3, LLMs can perform generating under the guidance of exemplars with in-context learning (Dong et al., 2023). By providing a relevant exemplar in the input, that is, one-shot prompting (Brown et al., 2020), LLMs can perform the CMG task with reference to the provided exemplar. The more relevant the provided exemplar is to the query diff, the more valuable it is as a reference, and LLMs are expected to generate better commit messages.
@@ 4. Experimental Setup
In this section, we first propose four research questions that reflect the basis of our experimental path. We then introduce the dataset used in the experiment, the selected models, the baselines for comparison, the metrics used for evaluating results, and the implementation details for reference and reproduction.
@@ 4.1. Research Questions
The overall goal of this research is to validate the effectiveness of REACT. To achieve this goal, we conducted experiments aimed at investigating the following research questions (RQs):
@@ 4.2. Dataset
We employ a well-established dataset (Xu et al., 2019) in this study, which has been widely used in previous CMG works (Hoang et al., 2020; Xu et al., 2019; Dong et al., 2022; Jiang et al., 2017; Liu et al., 2018; Wang et al., 2022). The dataset is obtained from the top 1,000 popular Java repositories in GitHub, and a total of 90,661 pairs of code diff and corresponding commit messages are collected. Following the previous work (Xu et al., 2019), we divide the dataset into training, validation, and testing sets in the ratio shown in Table 1. In addition to using the original dataset to evaluate the performance of directly applying CLMs to the CMG task in RQ1, we also construct an augmented dataset. The augmented dataset is derived from the original dataset through a retrieval augmentation process described in Section 3.2 and Section 3.3. Thus, for each query diff in the original dataset, a relevant diff-message pair is retrieved and saved. The augmented dataset is used for assessing the effectiveness of REACT in RQ2–RQ4. Table 1 presents the statistics of both the original and augmented datasets, in which “Count” refers to the number of entries in each dataset split. We can see that since each entry of the augmented dataset includes two diff-message pairs (query and retrieved pair), the number of tokens nearly doubles. All the models in the experiments are evaluated on the testing set. The training and validation sets are used in the fine-tuning stage of PLMs.

### v3

@@ Abstract.
Commit message is one of the most important textual information in software development and maintenance. However, it is time-consuming to write commit messages manually. Commit Message Generation (CMG) has become a research hotspot. Recently, several pre-trained language models (PLMs) and large language models (LLMs) with code capabilities have been introduced, demonstrating impressive performance on code-related tasks. Meanwhile, prior studies have explored the utilization of retrieval techniques for CMG, but it is still unclear what effects would emerge from combining advanced retrieval techniques with various generation models. This paper proposed REACT, a REtrieval-Augmented framework for CommiT message generation. It integrates advanced retrieval techniques with different PLMs and LLMs, to enhance the performance of these models on the CMG task. Specifically, a hybrid retriever is designed and used to retrieve the most relevant code diff and commit message pair as an exemplar. Then, the retrieved pair is utilized to guide and enhance the CMG task by PLMs and LLMs through fine-tuning and in-context learning. The experimental results show that REACT significantly enhances these models’ performance on the CMG task, improving the BLEU score of CodeT5 by up to 55%, boosting Llama 3’s BLEU score by 102%, and substantially surpassing all baselines.
@@ 1. Introduction
In this context, Commit Message Generation (CMG) has become a hot topic in automated software engineering and has garnered attention from this research community. CMG task aims to take the differences between two versions of code as input, typically in the form of a code diff file .diff generated by Git, aka code change, and then generates the corresponding commit message. Initially, rule-based approaches were used in CMG (Shen et al., 2016; Linares-Vásquez et al., 2015; Buse and Weimer, 2010), where predefined rules or templates were utilized for generation. Some retrieval-based approaches leverage information retrieval (IR) techniques to suggest commit messages from similar code diff (Liu et al., 2018; Hoang et al., 2020). With the rapid development of deep learning, many learning-based methods have recently emerged (Jiang et al., 2017; Loyola et al., 2017; Xu et al., 2019; Dong et al., 2022; Liu et al., 2019). These methods treat CMG as a neural machine translation task, in which code diff is the input of neural network model to generate commit messages as output. In addition, there are some hybrid approaches (Liu et al., 2022; Wang et al., 2022; Shi et al., 2022) that leverage both IR and deep learning techniques.
More recently, various language models possessing code capabilities have been proposed, including code-specific pre-trained language models such as CodeT5 (Wang et al., 2021) and general large language models, notably ChatGPT [^1] https://chatgpt.com/, which have attracted the vast attention of the research community. In this paper, we refer to these code-capable language models as Code Language Models (CLMs). Different from traditional models, CLMs can be adapted for downstream code-related tasks by merely fine-tuning or prompting (Wei et al., 2021), rather than training a model from initial weights. Such approaches not only conserve resources and training time but also can achieve better performance. Researchers have explored applying them to the CMG task (Zhang et al., 2024c, a) and got some promising prospects.
@@
In this paper, we propose REACT as a REtrieval-Augmented framework for CommiT message generation. REACT comprises three phases, i.e., Retrieve, Augment, and Generate, to generate commit messages for given code diff. Retrieve: We design a hybrid retriever to retrieve the most relevant diff and its corresponding commit message from a comprehensive and high-quality source database. The retrieved diff-message pair will be used to guide the subsequent generation of commit message. Augment: The query diff and retrieved diff-message pair need to be combined as the input of the models, that is, input augmentation. The augmenter concatenates them with pre-defined special tokens or fills them into a prompt template before passing the result to the generator for the next phase. Generate: The generator receives augmented inputs and generates a commit message corresponding to the query diff under the guidance of the retrieved pair. We use various CLMs as the backbone of the generator instead of training a model from scratch. This can fully leverage the prior knowledge in CLMs and generate messages more effectively. Some generators with PLM backbones require training to learn how to leverage the additional information provided by the retrieved pair to enhance commit message generation. If LLM is used as the generator, the enhancement can be achieved through in-context learning (Dong et al., 2023) by embedding the retrieved pairs into the input prompt, eliminating the need for training. The final commit message generated through three stages is enhanced by the additional information or similar patterns provided by the exemplar.
@@ Results.
We designed and conducted comprehensive experiments to evaluate REACT on a widely-used CMG dataset. Before assessing the enhancement effects of REACT, we first reported the performance of directly using CLMs to generate commit messages. The experimental results showed that without integrating the REACT framework, directly generating commit messages with PLMs yielded remarkable results, surpassing all baselines comprehensively. Specifically, CodeT5’s BLEU score was 21% higher than the best baseline method, demonstrating the superiority of PLMs. Although LLMs showed mediocre scores on some metrics, Llama 3’s METEOR score still surpassed all baselines without additional training. When integrating CLMs into REACT, the model’s performance had broad and significant improvements. CodeT5’s BLEU score increased by 55% compared to direct application, and Llama 3’s BLEU score increased by 102%. The results established the effectiveness and wide applicability of the REACT framework, setting a new SOTA for the CMG task. Furthermore, we conducted a within-project case study on Electron, a popular open-source project from GitHub, where the exemplars were retrieved solely from the project’s historical commits. The results indicated that guided by exemplars, the generation process adheres well to the project’s commit message writing conventions, thereby significantly improving metric scores. This finding provides valuable insights into the application of REACT in real-world within-project CMG scenarios.
@@
  • An effective retrieval-augmented framework for CMG. We propose REACT, a retrieval-augmented framework for enhancing commit message generation. REACT can effectively integrate advanced retrieval techniques with various models for better CMG.
  • Extensive experiments are designed to evaluate the performance of various CLMs on the CMG task. To the best of our knowledge, this work is the first to comprehensively evaluate the performance of various CLMs on the CMG task, including PLMs (e.g., CodeT5, UniXCoder, etc.) and LLMs (e.g., ChatGPT, Llama 3, Gemma). Experimental results show that, benefiting from the code understanding ability acquired during pre-training, CLMs achieved impressive performance on CMG with simple fine-tuning or prompting with basic instruction, surpassing the baselines by a large margin. Specifically, CodeT5’s BLEU score was 21% higher than the best baseline method RACE.
@@
The remainder of this paper is structured as follows: Section 2 covers related work on CMG and CLMs. Section 3 introduces our REACT framework, detailing its three phases. Section 4 describes the experimental setup, including the dataset and model selection, while Section 5 presents results and analysis. Finally, Section 6 and Section 7 discuss the implications and threats to validity respectively, followed by the conclusions in Section 8. The replication package is available online (Zhang et al., 2024b)
@@ 3. Methodology
This section presents the methodology of our proposed three-phrase framework, i.e., REACT. We first provide a brief overview of REACT. Then, the three phases of REACT are addressed in detail.
@@ 3.1. Overview
As shown in Fig. 2, REACT comprises three phases, i.e., Retrieve, Augment, and Generate, to generate the commit message for a specified code diff. In this paper, we refer to the code diff that awaits the generation of its corresponding commit message as the “query diff”. Firstly, REACT retrieves the most relevant diff-message pair from the source database using a hybrid retriever. Then, the query diff and the retrieved diff-message pair are sent together to the augmenter for augmentation. Finally, the augmented input is fed into the generator to generate a commit message under the guidance of the retrieved pair.
@@ 3.2. Phase I: Retrieve
Phase I intends to retrieve the most relevant diff-message pair for guidance. which is accomplished through the following two components:
@@ 3.2.1. Hybrid Retriever
To more accurately assess the similarity between two code diffs and retrieve the most relevant diff-message pair based on that similarity, we designed an advanced hybrid retriever that involves combining two similarity scores with weighted fusion. BM25 (Robertson and Zaragoza, 2009) is a relevance scoring algorithm commonly used in information retrieval and search engines. It measures the relevance of a document to a given query by considering factors such as term frequency, document length, and term saturation. BM25 treats query diff as a bag-of-words representation and computes lexical similarity scores between query diff and each of the candidates. We also employ a transformer encoder-based similarity scoring method. As an improved version of CodeT5 (Wang et al., 2021), CodeT5+ (Wang et al., 2023a) is a cutting-edge PLM for code with the architecture of transformer encoder-decoder. Through its pre-training process, it learned rich representations from code data. We extract the encoder component of CodeT5+ and then utilize it to transform the code diff text into a dense 256-dimensional vector embedding, which is generally considered to encapsulate the semantic information of the code diff. By calculating the cosine similarity score between two vectors, we can obtain the similarity score between two code diffs. Following the normalization of the scores to a common scale, we combine the scores obtained from these two methods with equal weights (1:1) in our experiment, and we refer to it as the hybrid score. After calculating the hybrid scores between the query diff and all diffs in the source database, the diff-message pair with the highest score is retrieved as the most relevant pair. The hybrid retriever is expected to be more robust and effective compared to single retrievers that rely on only one similarity scoring method.
@@ 3.2.2. Source Database
To ensure that each query diff can retrieve a relevant diff-message pair as an exemplar and that the retrieved exemplar is of high quality, containing sufficient information to guide generation, a comprehensive, high-quality source database is crucial. To avoid reinventing the wheel, we reuse the data from CommitBench (Schall et al., 2024), a recently proposed dataset, to construct the source database used in our work. CommitBench collects commits from 72,000 publicly available GitHub repositories, which is the largest collection scope among existing datasets, ensuring the comprehensiveness of the dataset. Regarding the quality of this dataset, the repositories collected by CommitBench are non-fork repositories, which have been used by at least one other project and have a reasonable number of stars. Additionally, they performed an overall filtering process, designing ten filtering rules to exclude worthless messages, including bot messages, too-short or too-long messages, etc. The resulting source database contains over 1.6 million diff-message pairs. During the pre-processing stage, we persisted the 256-dimensional vector embedding of code diffs, obtained through the encoder in the hybrid retriever, as a new column in the source database to avoid additional computational overhead.
@@ 3.3. Phase II: Augment
This phase aims at combining the query diff and the retrieved pair into specific formats to augment the input context. For ease of reference, we denote <query diff>, <retrieved diff>, <retrieved msg> as the query diff for generation, the retrieved code diff, and the corresponding retrieved commit message, respectively. Considering the differences between PLMs and LLMs, we design the following two input augmentation methods for both:
@@ 3.3.1. For PLMs
PLMs require a specific input format to serve as the generator in REACT. To explicitly distinguish between different input components, we define three special tokens, i.e., [QUERY], [DIFF], and [MSG]. The augmenter concatenates these special tokens with query diff and retrieved pair into the following form:
@@ 3.3.2. For LLMs
We manually construct a prompt template for LLMs, as Figure 3 shows, to guide commit message generation by referring to retrieved exemplar. The augmenter fills in the corresponding content into the prompt template and then inputs it into LLMs for generation.
@@ 3.4. Phase III: Generate
The generator receives the augmented input from Phase II and generates the final commit message under the guidance of the retrieved pair inserted in the input. In this paper, we adopt various models, including PLMs and LLMs, as the generators and conduct a comprehensive study. The specific model selection will be detailed in Section 4.3. In this subsection, we discuss the two types of generators used in REACT and what adjustments they need to make to perform generating.
@@ 3.4.2. LLMs
In contrast to PLMs, LLMs can serve as the generator to generate commit messages without training. It benefits from the powerful generalization ability of LLMs, which can complete the specified tasks with only prompting (Wei et al., 2021). Meanwhile, through the well-designed instructions as shown in Figure 3, LLMs can perform generating under the guidance of exemplars with in-context learning (Dong et al., 2023). By providing a relevant exemplar in the input, that is, one-shot prompting (Brown et al., 2020), LLMs can perform the CMG task with reference to the provided exemplar. The more relevant the provided exemplar is to the query diff, the more valuable it is as a reference, and LLMs are expected to generate better commit messages.
@@ 4. Experimental Setup
In this section, we first propose four research questions that reflect the basis of our experimental path. Then, we introduce the dataset used in the experiments, the selected models, the baselines for comparison, the metrics used for evaluating results, as well as the implementation details for reference and reproduction.
@@ 4.1. Research Questions
The overall goal of this research is to validate the effectiveness of REACT. To achieve this goal, we conducted experiments aimed at investigating the following five research questions (RQs):
@@
RQ5: Can REACT perform well in the within-project scenario? Considering a within-project scenario where commit messages typically follow a consistent style, can REACT adapt its generated commit messages to align with the conventions of the specific project, enhancing both the relevance and usability of the output? This RQ is designed to investigate REACT’s ability to generate project-specific commit messages by leveraging historical records from the same project.
@@ 4.2. Dataset
In this work, we employ a well-established dataset (Xu et al., 2019) that has been widely used in previous CMG works (Hoang et al., 2020; Xu et al., 2019; Dong et al., 2022; Jiang et al., 2017; Liu et al., 2018; Wang et al., 2022). The dataset contains a total of 90,661 pairs of code diff and corresponding commit messages collected from the top 1,000 popular Java repositories in GitHub. Following the previous work (Xu et al., 2019), the dataset is divided into training, validation, and testing sets in the ratio, as shown in Table 1. In addition to using the original dataset to evaluate the performance of directly applying CLMs to the CMG task in RQ1, we also construct an augmented dataset. The augmented dataset is derived from the original dataset through a retrieval augmentation process described in Section 3.2 and Section 3.3. Thus, for each query diff in the original dataset, a relevant diff-message pair is retrieved and saved. The augmented dataset is used for assessing the effectiveness of REACT in RQ2–RQ4. Table 1 presents the statistics of both the original and augmented datasets, in which “Count” refers to the number of entries in each dataset split. We find that since each entry of the augmented dataset includes two diff-message pairs (query and retrieved pair), the number of tokens nearly doubles. All the models in the experiments are evaluated on the testing set. The training and validation sets are used in the fine-tuning stage of PLMs.